What is the area of square in chapter perimeter and area classs 7 maths 
 

Grade:7

1 Answers

Vedant Borkar
2 years ago
Dear, student
 Whenever asked about area of square,we need length of any one of the side (as all sides are equal in square).
 
So, assuming length of each side to be "a"..
We get area of square by the formula :
 
    Area =a×a=a^2 
 
Hope,its helpful XD